### v1

# Inverse Problems with Diffusion Models: A MAP Estimation Perspective

Abstract
Inverse problems have many applications in science and engineering. In Computer vision, several image restoration tasks such as inpainting, deblurring, and super-resolution can be formally modeled as inverse problems. Recently, methods have been developed for solving inverse problems that only leverage a pre-trained unconditional diffusion model and do not require additional task-specific training. In such methods, however, the inherent intractability of determining the conditional score function during the reverse diffusion process poses a real challenge, leaving the methods to settle with an approximation instead, which affects their performance in practice. Here, we propose a MAP estimation framework to model the reverse conditional generation process of a continuous time diffusion model as an optimization process of the underlying MAP objective, whose gradient term is tractable. In theory, the proposed framework can be applied to solve general inverse problems using gradient-based optimization methods. However, given the highly non-convex nature of the loss objective, finding a perfect gradient-based optimization algorithm can be quite challenging, nevertheless, our framework offers several potential research directions. We use our proposed formulation and develop empirically effective algorithms for solving noiseless and noisy image inpainting tasks. We validate our proposed algorithms with extensive experiments across diverse mask settings.
1 Introduction
Inverse problems are ubiquitous in science and engineering with a wide range of downstream applications  [27, 2]. In Computer vision, several image restoration tasks such as inpainting, deblurring, super-resolution, and more, can be formally modeled as inverse problems. In an inverse problem, characterized by Eq. 1, $y\in\mathbb{R}^{m}$ is a (potentially noisy) observation of the original data $x\in\mathbb{R}^{n}$, and $\eta$ is a random variable denoting i.i.d noise, typically assumed to be Gaussian with a known variance i.e $\eta\sim\mathcal{N}(0,\sigma^{2}_{y}\mathbb{I})$, and the task is to infer the original data $x$ given the observation $y$. The function $\mathcal{A}:\mathbb{R}^{n}\to\mathbb{R}^{m}$ is known as the forward operator, and typically $n\gg m$, indicating that the observation $y\in\mathbb{R}^{m}$ corresponds to a severely degraded signal, from which one needs to recover the original signal $x\in\mathbb{R}^{n}$, which makes the task highly challenging. For linear inverse problems, $\mathcal{A}$ denotes a linear mapping and can be substituted with a matrix $H\in\mathbb{R}^{m\times n}$.
$y=\mathcal{A}(x)+\eta$ (1)
Several conventional approaches for solving inverse problems exist [1]. These include approaches based on functional-analytic, probabilistic (bayesian), data-driven methods, and more. Recently, Deep Learning (DL) based methods have been applied to solve inverse problems and have shown great success. In a bayesian framework, solving an inverse problem naturally corresponds to estimating the posterior $P(x|y)$. Typical DL-based approaches for solving inverse problems fall into two categories. 1. Methods that directly learn the posterior $P(x|y)$ via conditional generative models [13, 18], and 2. Methods that learn $P(x)$ via an unconditional generative model and use it to infer $P(x|y)$  [20, 5, 28, 11]. Methods of the former category require task-specific training, i.e. training with a dataset of pairs $(x,y)$, where the degradation $y$ is computed using $x$ and a task-specific forward operator $\mathcal{A}$. This limits the out-of-the-box applicability of the model to a different task (different forward operator). On the contrary, methods of the latter category train an unconditional generative model to learn $P(x)$, and this training is task-independent since it only needs a dataset of original data samples $x$. These methods then use the trained model for $P(x)$ and since $P(y|x)$ is tractable ( i.e from Eq. 1, $P(y|x)=\mathcal{N}(\mathcal{A}(x),\sigma^{2}_{y}\mathbb{I})$ ), utilizing the bayes rule, they infer the posterior $P(x|y)\propto P(y|x)P(x)$.
Several choices for Deep Generative Models (DGMs) exist, each with its advantages and disadvantages. There have been approaches using Generative Adversarial Networks (GAN) [8] and Normalizing Flow (NF) [16] based DGMs for solving inverse problems, with more recent methods focusing on Diffusion models [19, 9, 25], owing to their state-of-the-art performance in several vision-based generative tasks. In this work, we focus on methods that use a pre-trained unconditional diffusion model as the prior $P(x)$ and infer the posterior $P(x|y)$ for solving inverse problems. Sec. 2 provides some background on diffusion models and related works that use unconditional diffusion models for solving inverse problems and their inherent limitations. Later in  Sec. 3, we propose our Maximum A Posteriori (MAP) estimation framework for continuous-time diffusion models and discuss the practical implementation. In  Sec. 4, use our proposed framework to develop empirically effective algorithms for image inpainting, and in  Sec. 5, we validate our proposed method with extensive experiments on image inpainting across diverse mask settings. Finally, in  Secs. 6 and 7, we discuss concurrent works and conclude our findings.
2 Background
Given a dataset of samples $\mathcal{D}={\{x_{i}\}}_{i=1}^{N}$, where each $x_{i}$ is an i.i.d sample drawn from an unknown data distribution $P_{data}(x)$, a generative model learns to approximate $P_{data}$ from the samples in $\mathcal{D}$.
2.1 Diffusion Models
Diffusion models  [19, 9, 25] are a recent family of generative models. The methodology involves simulating a stochastic process $\{x(t)\}_{t=0}^{T}$ described by an SDE (Stochastic Differential Equation) such as Eq. 2, where $t\in[0,T]$ is a continuous time variable, $x(0)\sim P_{0}=P_{data}$ is the data distribution for which we have a dataset $\mathcal{D}$ of samples, $x(T)\sim P_{T}$ is a tractable prior distribution. The functions $f(.,t):\mathbb{R}^{n}\to\mathbb{R}^{n}$ and $g(.):\mathbb{R}\to\mathbb{R}$ are called the drift and diffusion coefficients of $x(t)$ respectively, and $dw$ denotes the standard Wiener process. Typically $f$ and $g$ are chosen in a way that yields a tractable prior $P_{T}$ which contains no information about $P_{0}$ (i.e. $P_{data}$).
$dx=f(x,t)dt+g(t)dw$ (2)
Eq. 2 also describes the ”forward process”, in which, starting from an initially clean data sample, a small amount of noise is progressively added at each step until it turns into a noisy sample of the prior distribution $P_{T}$. The backward/reverse process which transforms a noisy sample of $P_{T}$ into a clean sample of the data distribution is described by the corresponding reverse-SDE in Eq. 3
$dx=[f(x,t)-g(t)^{2}\nabla_{x}\log P_{t}(x)]dt+g(t)d\bar{w}$ (3)
$d\bar{w}$ denotes the standard Wiener process, and $t$ goes backward from $T$ to 00 with $dt$ being an infinitesimal negative time step. The term $\nabla_{x}\log P_{t}(x)$ is called the score function of the marginal distribution $P_{t}(x)$. If we know this score function for each marginal distribution i.e. for all $t$, then one could, in theory, simulate the reverse-SDE in Eq. 3 and generate samples from the data distribution. In general, the score function is not analytically tractable and is hard to estimate, however, one could train a time-indexed neural network model to learn the score function via score matching techniques [26, 24]. The trained score model $S_{\theta}(x,t)$ can then be substituted in place of $\nabla_{x_{t}}\log P_{t}(x)$ in Eq. 3, and the reverse-SDE can be solved using traditional SDE solvers [25].
$dx=[f(x,t)-\frac{1}{2}g(t)^{2}\nabla_{x}\log P_{t}(x)]dt$ (4)
For the stochastic process described by the SDE in Eq. 2, there exists a corresponding ODE (Ordinary Differential Equation) shown in Eq. 4 describing a deterministic process whose trajectories share the same marginal probability densities $\{P(x_{t})\}_{t=0}^{T}$ as those simulated by the SDE. This is referred to as the probability flow ODE in the literature [25]. So equivalently, one could also use ODE solvers to solve Eq. 4 in reverse time from $t=T$ until 00 to generate samples from the data distribution.
From hereon, unless mentioned otherwise, we assume the default choice for drift and diffusion coefficients as $f(x,t)=0$ and $g(t)=\sqrt{\frac{d\sigma^{2}(t)}{dt}}$, where $\sigma(t)$ is a monotonically increasing noise schedule from $t=0$ to $T$, with $\sigma(T)$ being very high. This choice of $f$ and $g$ results in a closed form perturbation kernel $P(x(t)|x(0))=\mathcal{N}(x(0),\sigma^{2}(t)-\sigma^{2}(0))$ and $P(x(T))\approx\mathcal{N}(0,\sigma^{2}(T))$. There can be multiple design choices for the noise schedule $\sigma(t)$, resulting in various formulations of diffusion models  [10].
2.2 Solving Inverse problems with Diffusion models
As described in Sec. 1, solving an inverse problem entails estimation of (or sampling from) the posterior $P(x|y)$, where $y$ is the noisy degradation of $x$ (Eq. 1). In the context of solving inverse problems using diffusion models, sampling from the posterior $P(x|y)$ involves conditioning the reverse diffusion process on $y$ which translates to solving the modified reverse-SDE in Eq. 5.
$dx=[f(x,t)-g(t)^{2}\nabla_{x}\log P_{t}(x|y)]dt+g(t)d\bar{w}$ (5)
Similar to methods of the first category (Sec. 1), which directly learn the posterior $P(x|y)$ as part of their training, it is possible to train a conditional diffusion model that learns the conditional score function directly. More specifically, one could learn $S_{\theta}(x,y,t)$ using conditional score matching objectives in place of the usual unconditional score function $S_{\theta}(x,t)$ in Sec. 2.1. In this paper, we focus on methods of the second category, which only leverage an unconditional diffusion model for $P(x)$, to infer $P(x|y)$. We modify the notations to denote $P_{t}(x)$ with $P(x_{t})$ and $x(t)$ with $x_{t}$ for convenience.
$\nabla_{x_{t}}\log P(x_{t}|y)=\nabla_{x_{t}}\log P(x_{t})+\nabla_{x_{t}}\log P(y|x_{t})$
[…]
Solving Eq. 5 involves estimating the conditional score function $\nabla_{x_{t}}\log P(x_{t}|y)$. The pre-trained unconditional diffusion model can be used to estimate $\nabla_{x_{t}}\log P(x_{t})$, however, the term $\nabla_{x_{t}}\log P(y|x_{t})$ becomes intractable, making it hard to estimate the conditional score (Eqs. 6 and 7). At its core, the intractability of $\nabla_{x_{t}}\log P(y|x_{t})$ arises from the fact that $P(x_{0}|x_{t})$ is intractable [20].
2.3 Related works
PGDM [20] approximates $P(x_{0}|x_{t})$ with a Gaussian distribution having mean $\hat{x_{t}}$ and variance ${r_{t}}^{2}$, where $\hat{x_{t}}=\mathbb{E}(x_{0}|x_{t})=x_{t}+\sigma^{2}(t)\nabla_{x_{t}}\log P(x_{t})$ (using Tweedie’s formula). The standard deviation $r_{t}$ is a hyperparameter, chosen proportionally to $\sigma(t)$. DPS [5] approximates $P(y|x_{t})$ with the point estimate $P(y|x_{0}=\hat{x_{t}})$ and has an almost similar formulation as PGDM, though the motivation is slightly different.  [3] also uses Gaussian approximation but further replace $r_{t}$ with the covariance matrix $Cov[x_{0}|x_{t}]=\sigma^{2}(t)\frac{\partial\hat{x_{t}}}{\partial x_{t}}$. Computing this matrix is expensive in practice, so they resort to diagonal and row-sum approximations of the matrix instead.  [15] proposes to find an optimal covariance matrix using learned covariances from the diffusion model. All those methods, however, still rely on making simplified approximations for $P(x_{0}|x_{t})$. This limits their performance in practice, given the complicated and multimodal nature of the true data distribution. Other lines of work  [11, 29, 7, 6] try to circumvent this term by projecting the intermediate $x_{t}$ onto the measurement subspace using heuristic approximations. Some other works  [6, 31, 4, 28] pose the problem as a MAP optimization with data and prior terms. Note that our proposed method is quite similar to the concurrent works  [4, 28] from a practical perspective. We discuss the similarities and differences later in Sec. 6.
3 Our Methodology
3.1 Background: Consistency Models
Consider the probability flow ODE described in Eq. 4. The solution trajectories of this ODE are smooth, and maps the samples on the data manifold to pure noise. In  [22], a consistency model is defined as the function that map any point on the PF ODE trajectory to its corresponding origin (initial point on the data manifold). There exist efficient methodologies [21] to train these consistency models in practice. Please refer to  [22] for a detailed description of consistency models.
3.2 Proposed MAP estimation framework
$x^{*}_{0}=\arg\max_{x_{0}}\ \log P(x_{0}|y)$ (8)
Eq. 8 refers to the usual MAP formulation for solving an inverse problem. Finding an optimal $x^{*}_{0}$ using gradient ascent involves the update step in Eq. 9.
$x^{k+1}_{0}=x^{k}_{0}+\lambda*\nabla_{x_{0}}\log P(x_{0}|y)$ (9)
The update step requires computing the gradient term $\nabla_{x_{0}}\log P(x_{0}|y)=\nabla_{x_{0}}\log P(y|x_{0})+\nabla_{x_{0}}\log P(x_{0})$. The former term is tractable since $P(y|x_{0})$ is Gaussian, and the latter is the score function evaluated at $x_{0}$ and can be replaced with $S_{\theta}(x_{0},0)$.
In practice, $S_{\theta}(x_{0},0)$ is only accurate when $x_{0}$ lies closer to the data manifold and is typically inaccurate for $x_{0}$ in low-likelihood regions outside the data manifold (refer to  [23]). This makes the score estimate inaccurate at the start (when $x_{0}$ is initialized randomly) and also during the gradient ascent since the intermediate $x^{k}_{0}$ are not constrained to lie on the data manifold. This issue is avoided when inverse problems are typically solved through the reverse diffusion process (Sec. 2.2) which drifts the noisy sample towards the data manifold using $S_{\theta}(x_{t},t)$ while simultaneously ensuring measurement consistency. But there the bottleneck is to estimate $\nabla_{x_{t}}\log P(y|x_{t})$ which is intractable, as discussed earlier.
Here, we present our proposed MAP formulation. Let $z\sim P(x_{T})=\mathcal{N}(0,\sigma^{2}(T)\mathbb{I})$, denote a purely noisy sample, and $\mathcal{M}$ denote the data Manifold. The PF ODE trajectory maps $z$ to a sample $x_{0}\in\mathcal{M}$, given by $x_{0}=f_{\theta}(z,T)$, where $f_{\theta}$ is the consistency model. It is also evident that, $\forall\ x_{0}\in\mathcal{M}$, $\exists\ z\sim P(x_{T})$ such that $x_{0}=f_{\theta}(z,T)$. Hence, the usual MAP formulation in Eq. 8 is equivalent to the proposed MAP formulation in Eqs. 10 and 11.
$z^{*}=\arg\max_{z}\ \log P(x_{0}=f_{\theta}(z,T)|y)$ (10)
$x^{*}_{0}=f_{\theta}(z^{*},T)$ (11)
With our proposed formulation, we update $z$ with gradient ascent steps as in Eq. 12 for finding $z^{*}$. The update step now requires computing the gradient term $\nabla_{z}\log P(f_{\theta}(z,T)|y)$, which can be reformulated as a vector-jacobian product (vjp) as shown in Eq. 13.
The vector in the vjp from Eq. 13, is the gradient term $\nabla_{x_{0}}\log P(x_{0}|y)$ evaluated at $x_{0}=f_{\theta}(z,T)$ which lies on the data manifold (by definition of consistency model), and can be accurately evaluated unlike the previous case.
$z^{k+1}=z^{k}+\lambda*\nabla_{z}\log P(f_{\theta}(z,T)|y)$ (12)
$\nabla_{z}\log P(f_{\theta}(z,T)|y)={\left(\frac{\partial f_{\theta}(z,T)}{\partial z}\right)}^{\intercal}\nabla_{x_{0}}\log P(x_{0}|y)\bigg{|}_{x_{0}=f_{\theta}(z,T)}$
Here we provide the high-level overview of a practical implementation using our MAP formulation. In practice, even a consistency model $f_{\theta}$ can benefit from multi-step sampling  [22]. Therefore we propose a multi-step gradient ascent scheme called MAP-Gradient-Ascent (MAP-GA) as described in Algorithm 1. Note that $\tau$ refers to a time-step schedule with $T=\tau_{n}>\tau_{n-1}>..>\tau_{1}>\tau_{0}=0$ and $\sigma$ refers to the monotonically increasing noise schedule for $t\in[0,T]$, with $\sigma_{0}=0,$ and $\sigma_{T}=\infty$ (high value in practice), $y$ is the measurement, $f_{\theta}$ and $S_{\theta}$ are the consistency model and the score model respectively. $num\_iter$ denotes the number of gradient ascent iterations per time step, and $\lambda$ denotes the learning rate. Algorithm 1 can be applied to solve any inverse problem effectively, given that we know the optimal hyper-parameters, such as the time-step schedule $\tau$, the learning rate $\lambda$, learning rate schedule, $num\_steps$, etc. However, finding those in practice can be quite challenging.
input : $\tau=(\tau_{n},..\tau_{1},\tau_{0})$, $f_{\theta}$, $S_{\theta}$, y, num_iter, $\lambda$, $\sigma$
$z\sim\mathcal{N}(0,\sigma^{2}_{\tau_{n}}\mathbb{I})$
for $i\ in\ (n,n-1,..,1)$ do
       $t=\tau_{i}$
       for $j\ in\ (1,2,..,num\_iter)$ do
             $z=z+\lambda*\nabla_{z}\log P(f_{\theta}(z,t)|y)$
       end for
      $\hat{x}_{0}=f_{\theta}(z,t)$
       $z=\mathcal{N}(\hat{x}_{0},\sigma^{2}_{\tau_{i-1}}\mathbb{I})$
end for
output : z
Algorithm 1 MAP-GA (MAP-Gradient-Ascent)
3.3 Practical Implementation
In practice, to avoid numerical issues, instead of solving for $x^{*}_{0}$, we solve for $x^{*}_{\epsilon}=\arg\max_{x_{\epsilon}}\ \log P(x_{\epsilon}|y)$, for a small $\epsilon$ such that $\sigma_{\epsilon}\approx 0$. So we solve the MAP formulation in Eqs. 14 and 15. Since, $P(x_{\epsilon}|x_{0})=\mathcal{N}(x_{0},\sigma^{2}_{\epsilon}\mathbb{I})$, for very small values of $\sigma_{\epsilon}$, the distinction between $x_{0}$ and $x_{\epsilon}$ remain insignificant for all practical purposes. The consistency models in  [22] are also learned to map the points on PF ODE trajectory to the corresponding $x_{\epsilon}$ instead of $x_{0}$.
$z^{*}=\arg\max_{z}\ \log P(x_{\epsilon}=f_{\theta}(z,T)|y)$ (14)
$x^{*}_{0}\approx x^{*}_{\epsilon}=f_{\theta}(z^{*},T)$ (15)
Here, we describe in detail the computation of the gradient term $\nabla_{z}\log P(f_{\theta}(z,t)|y)$ in practice. From Eqs. 16 and 17, this requires the estimation of the gradient of log-likelihood i.e $\nabla_{x_{\epsilon}}\log P(y|x_{\epsilon})$ and the gradient of log-prior i.e $\nabla_{x_{\epsilon}}\log P(x_{\epsilon})$. The terms $P(y|x_{\epsilon})$ and $P(x_{\epsilon})$ are also referred to as the likelihood and the prior respectively.
$\nabla_{z}\log P(f_{\theta}(z,t)|y)={\left(\frac{\partial f_{\theta}(z,t)}{\partial z}\right)}^{\intercal}\nabla_{x_{\epsilon}}\log P(x_{\epsilon}|y)\bigg{|}_{x_{\epsilon}=f_{\theta}(z,t)}$
$\nabla_{x_{\epsilon}}\log P(x_{\epsilon}|y)\bigg{|}_{x_{\epsilon}=f_{\theta}(z,t)}=\biggl{\{}\nabla_{x_{\epsilon}}\log P(y|x_{\epsilon})+\nabla_{x_{\epsilon}}\log P(x_{\epsilon})\biggr{\}}\bigg{|}_{x_{\epsilon}=f_{\theta}(z,t)}$
Computing the gradient of log-likelihood
$P(y|x_{0})=\mathcal{N}(\mathcal{A}(x_{0}),\sigma^{2}_{y}\mathbb{I})$, and $P(x_{\epsilon}|x_{0})=\mathcal{N}(x_{0},\sigma^{2}_{\epsilon}\mathbb{I})$, given by Eq. 1 and the diffusion perturbation kernel respectively. Since $\sigma_{\epsilon}\approx 0$, $P(x_{0}|x_{\epsilon})$ can be approximated with $P(x_{\epsilon}|x_{0})$, and for a linear forward operator i.e $\mathcal{A}=H\in\mathbb{R}^{m\times n}$, we can derive $P(y|x_{\epsilon})=\mathcal{N}(Hx_{\epsilon},\sigma^{2}_{y}\mathbb{I}+\sigma^{2}_{\epsilon}HH^{\intercal})$. For non-linear $\mathcal{A}$, by linearizing it around $x_{\epsilon}$, similar approximations can be derived. As $\sigma_{\epsilon}\to 0$, these approximations tend to hold well. Given the tractable form of the likelihood term above, computing the gradient of the log-likelihood is apparent.
Computing the gradient of log-prior
The gradient of the log prior i.e $\nabla_{x_{\epsilon}}\log P(x_{\epsilon})$ is essentially the score function evaluated at $x_{\epsilon}$. Given a score function $S_{\theta}(x,t)$, learned by the unconditional diffusion model, $\nabla_{x_{\epsilon}}\log P(x_{\epsilon})=S_{\theta}(x_{\epsilon},\epsilon)$. Learning the score function is equivalent to learning a denoiser, and vice-versa [10]. Hereon, we denote the unconditional diffusion model as learning the denoiser $D_{\theta}(x,t)$, from which the score function can be computed as $\nabla_{x_{t}}\log P(x_{t})$ $=\frac{D_{\theta}(x_{t},t)-x_{t}}{\sigma^{2}_{t}}$.
4 Image Inpainting with MAP-GA
In Computer vision, image inpainting [30, 14] is a typical image restoration task. Given a masked image (and the corresponding binary mask), the goal here is to recover (reconstruct) the missing pixels of the masked image. Image inpainting can be modeled as a linear inverse problem. Let $x\in\mathbb{R}^{n}$ denote the original image, $y\in\mathbb{R}^{m}$ denote a masked image with only visible pixels ($m\leq n$) and $H\in\mathbb{R}^{m\times n}$ denotes a corresponding forward operator for a given mask. The inpainting problem is characterized by $y=Hx+\eta$, where $\eta\sim\mathcal{N}(0,\sigma^{2}_{y}\mathbb{I})$. Note that a given mask defines $H$ and for inpainting, $H$ has a defined structure, where the rows of $H$ are one-hot and are orthogonal. This also implies $HH^{\intercal}=\mathbb{I}_{m\times m}$.
input :  time schedule: $\tau=[\ \tau_{n},\tau_{n-1},..\tau_{1},\tau_{0}\ ]$,
noise schedule: $\sigma(.)$, Denoiser: $D_{\theta},$
consistency model: $C_{\theta}$,
measurement: $y$, learning rate: $\lambda,$
num gradient ascent iter: num_iter,
boolean: use_prior (default True),
Inpainting matrix: H,
Measurement noise: $\sigma_{y}$
(Note: $HH^{\intercal}=\mathbb{I}_{m\times m}$, $\tau_{0}=\epsilon$, $\tau_{n}=T$),
$z\sim\mathcal{N}(0,\sigma^{2}_{\tau_{n}}\mathbb{I})$
[…]
       $t=\tau_{i}$
[…]
             $\hat{x}_{\epsilon}=C_{\theta}(z,t)$
             $grad_{likelihood}=\frac{H^{\intercal}(y-H\hat{x}_{\epsilon})}{\sigma^{2}_{y}+\sigma^{2}_{\epsilon}}$
             if use_prior then
                  $grad_{prior}=\frac{D_{\theta}(\hat{x}_{\epsilon},\epsilon)-\hat{x}_{\epsilon}}{\sigma^{2}_{\epsilon}}$
             end if
            else
                  $grad_{prior}=0$
             end if
            $grad_{posterior}=grad_{likelihood}+grad_{prior}$
             $grad=\left(\frac{\partial C_{\theta}(z,t)}{\partial z}\right)^{\intercal}grad_{posterior}$
             $z=z+\lambda*grad$
       end for
      $\hat{x}_{\epsilon}=C_{\theta}(z,t)$
       $z=\mathcal{N}(\hat{x}_{\epsilon},\sigma^{2}_{\tau_{i-1}}-\sigma^{2}_{\tau_{0}}\mathbb{I})$
end for
output : z
Algorithm 2 MAP-GA (Algorithm 1 expanded for Inpainting)
In this work, to show the effectiveness of our proposed MAP framework in practice, we consider the task of image inpainting with diverse mask settings. We expand Algorithm 1 for the task of image inpainting, and include all the specific details in Algorithm 2. The core quantity in the algorithm that needs to be evaluated is $\nabla_{z}\log P(f_{\theta}(z,t)|y)$, which involves the estimation of gradients of the log-likelihood and the log-prior terms (Eqs. 16 and 17). In the algorithm, we make a choice (indicated by the $use\_prior$ keyword) of retaining or dropping the gradient of the log-prior. Dropping the prior term implies a choice of uniform prior, and the algorithm now optimizes for the maximum likelihood estimate instead of the MAP estimate, by considering any sample (which is consistent with our measurement) on the data manifold to be equally good. This is also the setting considered in the concurrent works  [4, 28].
Note that the algorithm makes use of both the consistency model ($C_{\theta}$) and the denoiser ($D_{\theta}$), but this makes the method more demanding compared to other methods that only use the denoiser. For that, we make an argument as follows. We use the pre-trained denoiser and the consistency model from  [10] and  [22] respectively, with noise schedule $\sigma(t)=t$ and $t\in[\epsilon,T]$, where $\epsilon=0.002,T=80$. Consider the PF ODE for the above setting, as follows.
$dx$ $=-\sigma(t)\dot{\sigma}(t)\nabla_{x_{t}}\log P(x_{t})dt$
$=-\sigma(t)\dot{\sigma}(t)\frac{D_{\theta}(x,t)-x}{\sigma^{2}(t)}dt$
$=-\dot{\sigma}(t)\frac{D_{\theta}(x,t)-x}{\sigma(t)}dt$
$=-\frac{D_{\theta}(x,t)-x}{t}dt$
This PF ODE determines the trajectory and solving the trajectory origin $x_{0}$ involves solving the ODE above. Note that this $x_{0}$ is what the consistency model $C_{\theta}$ is trained to predict. As a rough approximation, solving the ODE with a backward Euler discretization step from $t$ to 00, which essentially assumes the trajectory curves are linear, i.e. the jacobian $\frac{dx}{dt}$ is constant for the interval $[0,t]$, and this gives $C_{\theta}(x_{t},t)=x_{0}\approx D_{\theta}(x_{t},t)$, with the approximation getting more accurate as the trajectory curve gets more linear. In an empirical setting,  [10] also observes trajectories become more linear when $\sigma(t)\to 0$. While more analysis on this is still due, it motivates us to look at the denoiser as a proxy of the consistency model, which gradually becomes more and more accurate as $\sigma(t)\to 0$. Hence, we also consider settings that replace the consistency model with the denoiser in our algorithm.
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Figure 1: From left to right: original image, and mask settings: box50, half, expand, box25, sr2x, altlines
input :  noise schedule: $\sigma(.)$, Denoiser: $D_{\theta},$
consistency model: $C_{\theta}$,
measurement: $y$, learning rate: $\lambda,$
num gradient ascent iter: num_iter,
Inpainting matrix: H,
Measurement noise: $\sigma_{y}$
$\tau^{map-ga}=[\tau^{map}_{n},\tau^{map}_{n-1},..\tau^{map}_{1},\tau^{map}_{0}]$,
$\tau^{pgdm}=[\tau^{pgdm}_{m},\tau^{pgdm}_{m-1},..\tau^{pgdm}_{1},\tau^{pgdm}_{0}]$,
(Note: $HH^{\intercal}=\mathbb{I}_{m\times m}$)
(Note: $\sigma_{\tau^{map}_{0}}=\sigma_{y}$ and $\tau^{map}_{n}=T$)
(Note: $\sigma_{\tau^{pgdm}_{m}}=\sigma_{y}$ and $\tau^{pgdm}_{0}=\epsilon$),
=======MAP-GA=======
$z\sim\mathcal{N}(0,\sigma^{2}_{\tau^{map}_{n}}\mathbb{I})$
[…]
       $t=\tau^{map}_{i}$
[…]
             $\hat{x}_{\epsilon}=C_{\theta}(z,t)$
             $grad_{likelihood}=\frac{H^{\intercal}(y-H\hat{x}_{\epsilon})}{\sigma^{2}_{y}+\sigma^{2}_{\epsilon}}$
             $grad_{posterior}=grad_{likelihood}$
             $grad=\left(\frac{\partial C_{\theta}(z,t)}{\partial z}\right)^{\intercal}grad_{posterior}$
             $z=z+\lambda*grad$
       end for
      $\hat{x}_{\epsilon}=C_{\theta}(z,t)$
       $z=\mathcal{N}(\hat{x}_{\epsilon},\sigma^{2}_{\tau^{map}_{i-1}}-\sigma^{2}_{\tau^{map}_{0}}\mathbb{I})$
end for
========PGDM========
$x_{\tau^{pgdm}_{m}}=z$
[…]
       $t=\tau^{pgdm}_{i}$
       $\hat{x}_{t}=D_{\theta}(z,t)$
       $\mu_{t}=\hat{x}_{t}+\sigma^{2}_{t}*\nabla_{x_{t}}\log P(y|x_{t})$pgdm update*
       $x_{\tau^{pgdm}_{i-1}}=\mathcal{N}(\mu_{t},\sigma^{2}_{\tau^{pgdm}_{i-1}}-\sigma^{2}_{\tau^{pgdm}_{0}}\mathbb{I})$
end for
output : $x_{\tau^{pgdm}_{0}}$
Algorithm 3 MAP-GA-PGDM for Inpainting
When using Algorithm 2 for noisy image inpainting ($\sigma_{y}>0$) in practice, we observe that it requires careful tuning of the learning rate and other hyperparameters. To avoid such additional and sensitive hyperparameters, we present Algorithm 3 for noisy inpainting, based on our empirical observations in Tab. 2. The algorithm requires no additional hyperparameters and works well in practice. The idea behind Algorithm 3, is to find a good approximate solution using MAP-GA at diffusion time $t=\tau$ where $\sigma_{\tau}=\sigma_{y}$, and then use it as an initialization for PGDM [20] for $\sigma_{t}<\sigma_{y}$. More specifically, we use MAP-GA until $\sigma_{t}=\sigma_{y}$, to estimate $x_{\sigma_{y}}$ and later use this as initialization to PGDM for $\sigma_{t}<\sigma_{y}$. (Note that we could use $t$ and $\sigma_{t}$ interchangeably in our notations since our noise schedule is $\sigma(t)=t$). We do not use the prior term for the MAP-GA part in Algorithm 3, as we find it more effective.
5 Experiments
In all our experiments, we use the pre-trained class-conditional denoiser from  [10] and the consistency model from  [22] with their default settings. Inspired by  [12], we evaluate the inpainting performance across six different mask settings that denote varying levels of degradation. Examples of our mask settings can be seen in Fig. 1. The mask settings box50 and box25 indicate a square crop at the center of the image, with the crop width equal to 50% and 25% of the image width respectively. In half, we mask out the right half of the image, the mask expand is the complement of box25. sr2x denotes a 2x super-resolution mask, and altlines masks out alternate rows of pixels. We evaluate the inpainted images using FID and LPIPS metrics.
In Tab. 1, we evaluate the performance of MAP-GA (Algorithm 2) for noiseless image inpainting (i.e. $\sigma_{y}=0$) on Imagenet [17] validation set with 64x64 resolution having 1k images. In the experiments, the setting MAP-GA denotes the default Algorithm 2, MAP-GA(NP) denotes MAP-GA with no prior, MAP-GA(D) denotes MAP-GA with denoiser replacing the consistency model, and MAP-GA(D,NP) denotes MAP-GA with no prior and the denoiser replacing the consistency model. The results from Tab. 1 show MAP-GA and variants outperforming PGDM by a significant margin. Note that MAP-GA uses gradient ascent (first-order gradient-based optimization algorithm) to optimize the underlying MAP objective which is typically highly non-convex and is not guaranteed to find the global optima. However, with an initialization, closer to the global optima, MAP-GA could converge.
box50 half expand box25 sr2x altlines
[…]
MAP-GA(D,NP) 32.312 0.096 37.986 0.157 $93.068$ $0.419$ 11.019 0.019 30.072 0.034 18.282 0.015
MAP-GA(D) $35.761$ $0.100$ $43.315$ $0.164$ $106.32$ $0.430$ $11.169$ $0.019$ $33.188$ $0.036$ $19.103$ $0.016$
MAP-GA(NP) $34.944$ $0.113$ $39.243$ $0.162$ 69.004 0.388 $12.818$ $0.027$ $30.303$ $0.035$ $20.321$ $0.018$
MAP-GA $36.733$ $0.113$ $41.151$ $0.164$ $87.952$ $0.400$ $12.836$ $0.027$ $33.752$ $0.036$ $20.895$ $0.018$
PGDM [20] $49.370$ $0.151$ $54.261$ $0.245$ $127.95$ $0.479$ $14.255$ $0.021$ $38.433$ $0.046$ $20.446$ $0.019$
Table 1: Noiseless inpainting on ImageNet64 1k validation set using MAP-GA and variants. The setting MAP-GA denotes the default Algorithm 2, MAP-GA(NP) denotes MAP-GA with no prior, MAP-GA(D) denotes MAP-GA with denoiser replacing the consistency model, and MAP-GA(D,NP) denotes MAP-GA with no prior and the denoiser replacing the consistency model.
box50 half expand box25 sr2x altlines
[…]
MAP-GA(D,NP) $26.661$ $0.043$ $30.621$ $0.064$ $73.510$ $0.155$ 9.165 $0.010$ $29.878$ $0.027$ 16.971 0.012
MAP-GA(D) $34.689$ $0.052$ $41.098$ $0.083$ $90.768$ $0.175$ $10.891$ $0.011$ $34.334$ $0.031$ $18.718$ $0.013$
MAP-GA(NP) 25.205 0.043 27.208 0.047 42.671 0.075 $10.936$ $0.017$ 27.496 0.023 $18.732$ $0.013$
MAP-GA $29.508$ $0.045$ $33.653$ $0.050$ $73.318$ $0.106$ $11.257$ $0.016$ $34.364$ $0.026$ $20.041$ $0.013$
PGDM [20] $31.952$ $0.056$ $36.639$ $0.082$ $71.352$ $0.138$ $10.886$ 0.009 $33.199$ $0.041$ $19.551$ $0.018$
Table 2: Noiseless inpainting on ImageNet64 1k validation set. Using the ground truth image ($x_{0}$) for the measurement $y$, we create a sample at $t=0.5$ via ($x_{0.5}=x_{0}+0.5*\eta$, where, $\eta\sim\mathcal{N}(0,\mathbb{I}))$ and initialize Algorithm 2 with $z=x_{0.5}$, and $\tau_{n}=0.5$
To corroborate this, we design the following toy experiment. We consider the same noiseless inpainting task from earlier, but, in Algorithm 2, we set $\tau_{n}=\hat{\tau}\ll T$, instead of the default setting $\tau_{n}=T$ and we also initialize $z\sim\mathcal{N}(x_{0},\sigma^{2}_{\hat{\tau}}\mathbb{I})$, where $x_{0}$ is the corresponding ground truth image for the measurement $y$. From Tab. 2, we observe significant improvements in the performance. (We use $\hat{\tau}=0.5$, note that $\sigma_{0.5}=0.5$, as we use the noise schedule $\sigma(t)=t$). This shows that MAP-GA is only limited by the choice of optimizer and reinforces the need for better optimization algorithms. While it is important to consider better design choices (for the schedules and hyperparameters), adaptive-gradient-based optimizers (such as momentum, adam), or higher-order methods, we leave this for future work as it requires a thorough analysis.
In Tab. 3, we report the results on noisy inpainting using Algorithm 3 on Imagenet 1k validation set with 64x64 resolution. $\sigma_{y}$ indicates the measurement noise and the setting MAP-GA-PGDM denote the default Algorithm 3, MAP-GA-PGDM(D) denote MAP-GA-PGDM with denoiser replacing the consistency model. In Algorithm 3, and all its variants, we fix $m=20$ time steps for PGDM. Even with a high level of measurement noise, MAP-GA-PGDM shows promising improvements over PGDM.
In all our experiments, each of our methods has been given a budget of 1000 steps, and we run ablations for (num_time_steps, num_iter) from the set {(20,50),(50,20),(100,10),(200,5),(250,4),(500,2),(1000,1)}. We follow a time-step schedule similar to  [10], with their default settings. The best results for all our methods resulted from the ablation (20,50) denoting 20 time-steps with 50 gradient ascent iterations. For PGDM [20], each setting from our results table is run with different choices of num_steps from {20,50,100,200,250,500,1000}, and most of the runs have their peak performance at $500$ steps. The performance reported in the tables is the best result for each method over the ablations for (num_time_steps, num_iter). The choice of learning rate in all our experiments was $\sigma^{2}_{y}+\sigma^{2}_{\epsilon}$.
box50 half expand box25 sr2x altlines
[…]
MAP-GA-PGDM(D) $0.05$ 56.173 0.125 $62.026$ $0.193$ $108.16$ $0.438$ 38.725 0.039 $57.41$ $0.080$ $46.046$ $0.046$
MAP-GA-PGDM $0.05$ $58.283$ $0.147$ 61.588 0.184 91.508 0.406 $44.667$ $0.085$ 57.308 0.073 45.265 0.042
PGDM [20] $0.05$ $77.824$ $0.175$ $80.248$ $0.257$ $135.99$ $0.495$ $53.136$ $0.049$ $86.289$ $0.126$ $66.203$ $0.066$
[…]
MAP-GA-PGDM(D) $0.1$ 72.543 0.166 $79.535$ $0.230$ $114.51$ $0.464$ 57.556 0.076 $76.733$ $0.145$ $65.154$ $0.096$
MAP-GA-PGDM $0.1$ $74.130$ $0.191$ 78.322 0.225 103.85 0.440 $65.720$ $0.161$ 76.248 0.134 63.134 0.089
PGDM [20] $0.1$ $96.485$ $0.216$ $99.170$ $0.286$ $145.40$ $0.519$ $78.620$ $0.100$ $109.47$ $0.231$ $90.925$ $0.138$
[…]
MAP-GA-PGDM(D) $0.2$ $91.112$ 0.243 $99.432$ $0.297$ $119.45$ $0.496$ 79.329 0.154 $110.48$ $0.260$ $87.727$ $0.188$
MAP-GA-PGDM $0.2$ 91.087 $0.252$ 97.531 0.296 116.78 0.491 $80.647$ $0.177$ 103.66 0.248 86.174 0.186
PGDM [20] $0.2$ $116.81$ $0.292$ $117.03$ $0.344$ $159.17$ $0.562$ $103.43$ $0.193$ $132.89$ $0.366$ $114.54$ $0.255$
Table 3: Noisy inpainting on ImageNet64 1k validation set. $\sigma_{y}$ denotes the measurement noise. The setting MAP-GA-PGDM denotes the default Algorithm 3, MAP-GA-PGDM(D) denote MAP-GA-PGDM with denoiser replacing the consistency model.
6 Discussion
The algorithms presented in this paper are similar to the concurrent works ZSIR [4], and DMPlug [28] from a practical perspective. However, our proposed MAP formulation has a strong theoretical motivation and connects the probability flow ODE and the consistency model with the MAP optimization for solving inverse problems. While we compare only against PGDM [20], it is the most competitive baseline and can also be seen from the experiments in DMPlug [28]. In this work, we analyze our method extensively, on image inpainting with various degradation masks. However, the experiments from DMPlug [28] further complement and validate our work. We show that the gradient of the log-prior is tractable, making the gradient of log-posterior tractable. However,  [4, 28] considers an arbitrary choice of uniform prior and ignores the prior term altogether. We show that MAP-GA is only limited by the choice of the optimizer in practice and our proposed framework has several directions for future research, such as improving optimization with hessian-based methods, more analysis on local optima, theory connecting the denoiser with the consistency model among others. Also, our method and algorithms are valid for general SDE-based Diffusion models and for more general inverse problems than the inpainting task we considered.
Figure 2: Noiseless inpainting example. From left to right: original image, masked image, MAP-GA(Algorithm 2), PGDM($\sigma_{y}=0$). From top to bottom: half mask, expand mask, box50 mask
Figure 3: Noisy inpainting example. From left to right: original image, masked image, MAP-GA-PGDM(Algorithm 3, $\sigma_{y}=0.1$), PGDM($\sigma_{y}=0.1$). From top to bottom: sr2x mask, box25 mask
7 Conclusion
In this paper, we proposed a novel MAP formulation for solving inverse problems using pre-trained unconditional diffusion models. Note that conditional generation is a core requirement in solving inverse problems. We connect the probability flow ODE and the consistency model with the optimization process for the MAP objective in tasks that involve conditional generation. We show that the gradient of the MAP objective is tractable, allowing the use of gradient-based optimization methods. To use our framework in practice, we propose an algorithm with a multi-step gradient ascent strategy for MAP optimization. Since we use gradient-ascent (a simple first-order gradient-based method), we observed that it requires extensive tuning of the learning rate and other hyperparameters due to the challenging optimization objective. To address this, we propose practically effective algorithms, for image inpainting using our framework (which can be extended to other inverse problems). We validate our framework with extensive experiments on image inpainting across diverse mask settings.
Acknowledgements
This work is part of the Marie Skłodowska-Curie Actions project MODELAIR, funded by the European Commission under the Horizon Europe program through grant agreement no. 101072559. The computations and the data handling were enabled by resources provided by the National Academic Infrastructure for Supercomputing in Sweden (NAISS), partially funded by the Swedish Research Council through grant agreement no. 2022-06725. Bharath thanks Sebastian Gerard and Heng Fang for their feedback on improving the presentation of this paper.