### v2

# Inverse Problems with Diffusion Models: A MAP Estimation Perspective

Abstract
Inverse problems have many applications in science and engineering. In Computer vision, several image restoration tasks such as inpainting, deblurring, and super-resolution can be formally modeled as inverse problems. Recently, methods have been developed for solving inverse problems that only leverage a pre-trained unconditional diffusion model and do not require additional task-specific training. In such methods, however, the inherent intractability of determining the conditional score function during the reverse diffusion process poses a real challenge, leaving the methods to settle with an approximation instead, which affects their performance in practice. Here, we propose a MAP estimation framework to model the reverse conditional generation process of a continuous time diffusion model as an optimization process of the underlying MAP objective, whose gradient term is tractable. In theory, the proposed framework can be applied to solve general inverse problems using gradient-based optimization methods. However, given the highly non-convex nature of the loss objective, finding a perfect gradient-based optimization algorithm can be quite challenging, nevertheless, our framework offers several potential research directions. We use our proposed formulation to develop empirically effective algorithms for image restoration. We validate our proposed algorithms with extensive experiments over multiple datasets across several restoration tasks.
1 Introduction
Inverse problems are ubiquitous in science and engineering with a wide range of downstream applications  [28, 2]. In Computer vision, several image restoration tasks such as inpainting, deblurring, super-resolution, and more, can be formally modeled as inverse problems. In an inverse problem, characterized by Eq. 1, $y\in\mathbb{R}^{m}$ is a (potentially noisy) observation of the original data $x\in\mathbb{R}^{n}$, and $\eta$ is a random variable denoting i.i.d. noise, typically assumed to be Gaussian with a known variance i.e $\eta\sim\mathcal{N}(0,\sigma^{2}_{y}\mathbb{I})$, and the task is to infer the original data $x$ given the observation $y$. The function $\mathcal{A}:\mathbb{R}^{n}\to\mathbb{R}^{m}$ is known as the forward operator, and typically $n\gg m$, indicating that the observation $y\in\mathbb{R}^{m}$ corresponds to a severely degraded signal, from which one needs to recover the original signal $x\in\mathbb{R}^{n}$, which makes the task highly challenging. For linear inverse problems, $\mathcal{A}$ denotes a linear mapping and can be substituted with a matrix $H\in\mathbb{R}^{m\times n}$.
$y=\mathcal{A}(x)+\eta$ (1)
Several conventional approaches for solving inverse problems exist [1]. These include approaches based on functional-analytic, probabilistic, data-driven methods, and more. Recently, Deep Learning (DL) based methods have been applied to solve inverse problems and have shown great success. In a Bayesian framework, solving an inverse problem naturally corresponds to estimating the posterior $P(x|y)$. Typical DL-based approaches for solving inverse problems fall into two categories. 1. Methods that directly learn the posterior $P(x|y)$ via conditional generative models [14, 19], and 2. Methods that learn $P(x)$ via an unconditional generative model and use it to infer $P(x|y)$  [21, 5, 29, 12]. Methods of the former category require task-specific training, i.e. training with a dataset of pairs $(x,y)$, where the degradation $y$ is computed using $x$ and a task-specific forward operator $\mathcal{A}$. This limits the out-of-the-box applicability of the model to a different task (different forward operator). On the contrary, methods of the latter category train an unconditional generative model to learn $P(x)$, and this training is task-independent since it only needs a dataset of original data samples $x$. These methods then use the trained model for $P(x)$ and since $P(y|x)$ is tractable (i.e. from Eq. 1, $P(y|x)=\mathcal{N}(\mathcal{A}(x),\sigma^{2}_{y}\mathbb{I})$), utilizing the Bayes rule, they infer the posterior $P(x|y)\propto P(y|x)P(x)$.
Several choices for Deep Generative Models (DGMs) exist, each with its advantages and disadvantages. There have been approaches using Generative Adversarial Networks (GAN) [9] and Normalizing Flow (NF) [17] based DGMs for solving inverse problems, with more recent methods focusing on Diffusion models [20, 10, 26], owing to their state-of-the-art performance in several vision-based generative tasks. This work focuses on methods that use a pre-trained unconditional diffusion model as the prior $P(x)$ and infer the posterior $P(x|y)$ for solving inverse problems. Sec. 2 provides some background on diffusion models and related works that use unconditional diffusion models for solving inverse problems and their inherent limitations. Later in  Sec. 3, we propose our Maximum A Posteriori (MAP) estimation framework for continuous-time diffusion models and discuss the practical implementation. Specifically, we propose a novel MAP formulation that employs a reparameterization based on consistency models [23] to model the reverse conditional diffusion process as MAP optimization. In  Sec. 4, we use our proposed framework to develop empirically effective algorithms for image restoration, and in  Sec. 5, we validate the algorithms with extensive experiments on deblurring, super-resolution, and image inpainting. In  Sec. 6, we present a brief discussion before concluding our findings in  Sec. 7.
2 Background
Given a dataset of samples $\mathcal{D}={\{x_{i}\}}_{i=1}^{N}$, where each $x_{i}$ is an i.i.d. sample drawn from an unknown data distribution $P_{data}(x)$, a generative model learns to approximate $P_{data}$ from the samples in $\mathcal{D}$.
2.1 Diffusion Models
Diffusion models  [20, 10, 26] are a recent family of generative models. The methodology involves simulating a stochastic process $\{x(t)\}_{t=0}^{T}$ described by a Stochastic Differential Equation (SDE) such as Eq. 2, where $t\in[0,T]$ is a continuous time variable, $x(0)\sim P_{0}=P_{data}$ is the data distribution for which we have a dataset $\mathcal{D}$ of samples, $x(T)\sim P_{T}$ is a tractable prior distribution. The functions $f(\cdot,t):\mathbb{R}^{n}\to\mathbb{R}^{n}$ and $g(\cdot):\mathbb{R}\to\mathbb{R}$ are called the drift and diffusion coefficients of $x(t)$ respectively, and $dw$ denotes the standard Wiener process. Typically $f$ and $g$ are chosen in a way that yields a tractable prior $P_{T}$ which contains no information about $P_{0}$ (i.e. $P_{data}$).
$dx=f(x,t)dt+g(t)dw$ (2)
Eq. 2 also describes the ”forward process”, in which, starting from an initially clean data sample, a small amount of noise is progressively added at each step until it turns into a noisy sample of the prior distribution $P_{T}$. The backward/reverse process which transforms a noisy sample of $P_{T}$ into a clean sample of the data distribution is described by the corresponding reverse-SDE in Eq. 3
$dx=[f(x,t)-g(t)^{2}\nabla_{x}\log P_{t}(x)]dt+g(t)d\bar{w}$ (3)
$d\bar{w}$ denotes the standard Wiener process when $t$ flows backwards from $T$ to 00 with $dt$ denoting an infinitesimal negative time step. The term $\nabla_{x}\log P_{t}(x)$ is called the score function of the marginal distribution $P_{t}(x)$. If we know this score function for each marginal distribution i.e. for all $t$, then one could solve the reverse-SDE in Eq. 3 and generate samples from the data distribution. In general, the score function is not analytically tractable and is hard to estimate, however, one could train a time-indexed neural network model to learn the score function via score matching techniques [27, 25]. The trained score model $S_{\theta}(x,t)$ can be substituted in place of $\nabla_{x}\log P_{t}(x)$ in Eq. 3, and the reverse-SDE can be solved using traditional SDE solvers [26].
$dx=[f(x,t)-\frac{1}{2}g(t)^{2}\nabla_{x}\log P_{t}(x)]dt$ (4)
For the stochastic process described by the SDE in Eq. 2, there exists a corresponding Ordinary Differential Equation (ODE) shown in Eq. 4, describing a deterministic process whose trajectories share the same marginal probability densities $\{P(x(t))\}_{t=0}^{T}$ as those simulated by the SDE. This is called the Probability Flow ODE (PF ODE) in the literature [26]. So equivalently, one could also use ODE solvers to solve Eq. 4 in reverse time from $t=T$ until 00 to generate samples from the data distribution.
Hereon, we assume the default choice for drift and diffusion coefficients as $f(x,t)=0$ and $g(t)=\sqrt{\frac{d\sigma^{2}(t)}{dt}}$, where $\sigma(t)$ is a monotonically increasing noise schedule from $t=0$ to $T$, with $\sigma(T)$ being very high. This choice of $f$ and $g$ results in a closed form perturbation kernel $P(x(t)|x(0))=\mathcal{N}(x(0),\sigma^{2}(t)-\sigma^{2}(0))$ and $P(x(T))\approx\mathcal{N}(0,\sigma^{2}(T))$. There can be multiple design choices for the noise schedule $\sigma(t)$, resulting in various formulations of diffusion models  [11].
2.2 Solving Inverse problems with Diffusion models
As described in Sec. 1, solving an inverse problem entails estimation of (or sampling from) the posterior $P(x|y)$, where $y$ is the noisy degradation of the original data sample $x$ (Eq. 1). In the context of solving inverse problems using diffusion models, sampling from the posterior $P(x(0)|y)$ involves conditioning the reverse diffusion process on $y$ which translates to solving the modified reverse-SDE in Eq. 5.
$dx=[f(x,t)-g(t)^{2}\nabla_{x}\log P_{t}(x|y)]dt+g(t)d\bar{w}$ (5)
Similar to methods of the first category (Sec. 1), which directly learn the posterior $P(x|y)$ as part of their training, it is possible to train a conditional diffusion model that learns the conditional score function directly. More specifically, one could learn $S_{\theta}(x,y,t)$ using conditional score matching objectives in place of the usual unconditional score function $S_{\theta}(x,t)$ in Sec. 2.1. In this work, we focus on methods of the second category, which only leverage an unconditional diffusion model for $P(x)$, to infer $P(x|y)$. We modify the notations to denote $x(t)$ with $x_{t}$, $\sigma(t)$ with $\sigma_{t}$, and $P_{t}(x)$ with $P(x_{t})$ for the sake of convenience and to be consistent with previous works [5, 21].
$\nabla_{x_{t}}\log P(x_{t}|y)=\nabla_{x_{t}}\log P(x_{t})+\nabla_{x_{t}}\log P(y|x_{t})$
[…]
Solving Eq. 5 involves estimating the conditional score function $\nabla_{x_{t}}\log P(x_{t}|y)$. The pre-trained unconditional diffusion model can be used to estimate $\nabla_{x_{t}}\log P(x_{t})$, however, the term $\nabla_{x_{t}}\log P(y|x_{t})$ becomes intractable (ref Eqs. 6 and 7). At its core, the intractability of $\nabla_{x_{t}}\log P(y|x_{t})$ arises from the fact that $P(x_{0}|x_{t})$ is intractable [21], and hence, the conditional score is hard to estimate while only leveraging the unconditional score.
2.3 Related works
PGDM [21] approximates $P(x_{0}|x_{t})$ with a Gaussian distribution having mean $\hat{x_{t}}$ and variance ${r_{t}}^{2}$, where $\hat{x_{t}}=\mathbb{E}(x_{0}|x_{t})=x_{t}+\sigma^{2}_{t}\nabla_{x_{t}}\log P(x_{t})$ (using Tweedie’s formula). The standard deviation $r_{t}$ is a hyperparameter, chosen proportionally to $\sigma_{t}$. DPS [5] approximates $P(y|x_{t})$ with the point estimate $P(y|x_{0}=\hat{x_{t}})$ and has an almost similar formulation as PGDM up to a constant factor, though the motivation is slightly different.  [3] also uses Gaussian approximation but further replace $r_{t}$ with the covariance matrix $Cov[x_{0}|x_{t}]=\sigma^{2}_{t}\frac{\partial\hat{x_{t}}}{\partial x_{t}}$. Computing this matrix is expensive in practice, so they resort to diagonal and row-sum approximations of the matrix instead.  [16] proposes to find an optimal covariance matrix using learned covariances from the diffusion model. All these methods, however, still make simplified approximations for $P(x_{0}|x_{t})$, which limits their performance in practice, given the complicated and multimodal nature of the true data distribution. Other lines of work  [12, 30, 7, 6] try to circumvent this term by projecting the intermediate $x_{t}$ onto the measurement subspace using heuristic approximations.
DiffPIR [33] poses the problem as MAP optimization with data and prior terms and utilizes the HQS [8] algorithm to solve a relaxed problem where the data and the prior terms can be optimized alternatively in a decoupled manner. Specifically, during each reverse diffusion step, this amounts to solving a relaxed MAP objective. DDS [6] uses a similar framework but employs subspace projection methods to solve the intermediate MAP objectives at each diffusion step. ZSIR [4] and DMPlug [29] also pose the problem as MAP optimization, where instead of optimizing for the original data $x_{0}$ directly, they reparameterize $x_{0}$ via the initial diffusion noise $x_{T}$ and solve for the optimal noise $x_{T}$ instead. From a theoretical perspective, it is unclear how this reparameterization should help. Also, both works ignore the prior term while focusing only on the data term. If the data distribution has full support i.e. $P(x_{0})\neq 0$ $\forall x_{0}\in\mathbb{R}^{n}$, ignoring the prior term, especially in case of a noisy measurement $y$ is theoretically unjustified. Our proposed method in this paper is similar to ZSIR and DMPlug from a practical perspective, however, our MAP formulation and algorithms are based on sound theoretical motivation that justifies the reparameterization based on PF ODE and provides new insights into modeling the conditional generation process as MAP optimization.
3 Our Methodology
3.1 Background: Consistency Models
Consider the PF ODE described in Eq. 4. The solution trajectories of this ODE are smooth, and map the samples on the data manifold to pure noise. In  [23], a consistency model is defined as the function that maps any point on the PF ODE trajectory to its corresponding origin (initial point on the data manifold). There exist efficient methodologies [22] to train these consistency models in practice. Please refer to  [23] for a detailed description.
3.2 Proposed MAP estimation framework
$x^{*}_{0}=\arg\max_{x_{0}}\ \log P(x_{0}|y)$ (8)
Eq. 8 refers to the usual MAP formulation for solving an inverse problem. Finding an optimal $x^{*}_{0}$ using gradient ascent involves the update step in Eq. 9, with $k$ denoting the $k^{th}$ iterate and $\lambda$ denoting the step size.
$x^{k+1}_{0}=x^{k}_{0}+\lambda*\nabla_{x_{0}}\log P(x_{0}|y)$ (9)
The update step requires computing the gradient term $\nabla_{x_{0}}\log P(x_{0}|y)=\nabla_{x_{0}}\log P(y|x_{0})+\nabla_{x_{0}}\log P(x_{0})$. The former term is tractable since $P(y|x_{0})$ is Gaussian, and the latter is the score function evaluated at $x_{0}$ and can be replaced with $S_{\theta}(x_{0},0)$. In practice, $S_{\theta}(x_{0},0)$ is only accurate when $x_{0}$ lies closer to the data manifold and is typically inaccurate for $x_{0}$ in low-likelihood regions outside the data manifold [24]. This makes the score estimate inaccurate in the beginning (when $x_{0}$ is initialized randomly) and during the gradient ascent updates, since the intermediate $x^{k}_{0}$ are not constrained to lie on the data manifold. This issue can be avoided when inverse problems are typically solved through the reverse diffusion process (Sec. 2.2) which drifts the noisy sample towards the data manifold using $S_{\theta}(x_{t},t)$ while simultaneously ensuring measurement consistency. But there the challenge is to estimate $\nabla_{x_{t}}\log P(y|x_{t})$ which is again intractable, as discussed earlier.
Here, we present our proposed MAP formulation. Let $z\sim P(x_{T})=\mathcal{N}(0,\sigma^{2}_{T}\mathbb{I})$, denote a purely noisy sample, and $\mathcal{M}$ denote the data Manifold. The PF ODE trajectory maps $z$ to a sample $x_{0}\in\mathcal{M}$, given by $x_{0}=f_{\theta}(z,T)$, where $f_{\theta}$ is the consistency model. It is also evident that, $\forall\ x_{0}\in\mathcal{M}$, $\exists\ z\sim P(x_{T})$ such that $x_{0}=f_{\theta}(z,T)$. Hence, the usual MAP formulation in Eq. 8 is equivalent to the proposed MAP formulation in Eqs. 10 and 11.
$z^{*}=\arg\max_{z}\ \log P(x_{0}=f_{\theta}(z,T)|y)$ (10)
$x^{*}_{0}=f_{\theta}(z^{*},T)$ (11)
With our proposed formulation, we update $z$ with gradient ascent steps as in Eq. 12 for finding $z^{*}$. The update step now requires computing the gradient term $\nabla_{z}\log P(f_{\theta}(z,T)|y)$, which can be reformulated as a vector-Jacobian product (vjp) as shown in Eq. 13. The vector in this vjp is the gradient term $\nabla_{x_{0}}\log P(x_{0}|y)$ evaluated at $x_{0}=f_{\theta}(z,T)$ which lies on the data manifold (by definition of consistency model) and can be accurately evaluated, unlike the previous case.
$z^{k+1}=z^{k}+\lambda*\nabla_{z}\log P(f_{\theta}(z,T)|y)$ (12)
$\nabla_{z}\log P(f_{\theta}(z,T)|y)={\left(\frac{\partial f_{\theta}(z,T)}{\partial z}\right)}^{\intercal}\nabla_{x_{0}}\log P(x_{0}|y)\bigg{|}_{x_{0}=f_{\theta}(z,T)}$
Here we provide a high-level overview of a practical implementation using our MAP formulation. In practice, even a consistency model $f_{\theta}$ can benefit from multi-step sampling  [23]. Therefore we propose a multi-step gradient ascent scheme called MAP-Gradient-Ascent (MAP-GA) as described in Algorithm 1. Note that $\tau$ refers to a time-step schedule with $T=\tau_{n}>\tau_{n-1}>..>\tau_{1}>\tau_{0}=0$ and $\sigma$ refers to the monotonically increasing noise schedule for $t\in[0,T]$, with $\sigma_{0}=0,$ and $\sigma_{T}=\infty$ (high value in practice), $y$ is the measurement, $f_{\theta}$ and $S_{\theta}$ are the consistency model and the score model respectively. $num\_iter$ denotes the number of gradient ascent iterations per time step, and $\lambda$ denotes the learning rate. Algorithm 1 can be applied to solve any inverse problem effectively, given that we know the optimal hyper-parameters, such as the time-step schedule $\tau$, the learning rate $\lambda$, learning rate schedule, $num\_steps$, etc. However, finding those in practice can be quite challenging.
input : $\tau=(\tau_{n},..\tau_{1},\tau_{0})$, $f_{\theta}$, $S_{\theta}$, y, num_iter, $\lambda$, $\sigma$
$z\sim\mathcal{N}(0,\sigma^{2}_{\tau_{n}}\mathbb{I})$
for $i\ in\ (n,n-1,..,1)$ do
       $t=\tau_{i}$
       for $j\ in\ (1,2,..,num\_iter)$ do
             $z=z+\lambda*\nabla_{z}\log P(f_{\theta}(z,t)|y)$
       end for
      $\hat{x}_{0}=f_{\theta}(z,t)$
       $z=\mathcal{N}(\hat{x}_{0},\sigma^{2}_{\tau_{i-1}}\mathbb{I})$
end for
output : z
Algorithm 1 MAP-GA (MAP-Gradient-Ascent)
3.3 Practical Implementation
In practice, to avoid numerical issues and to ensure that the gradient term $\nabla_{x_{0}}\log P(x_{0})$ exists, instead of solving for $x^{*}_{0}$, we solve for $x^{*}_{\epsilon}=\arg\max_{x_{\epsilon}}\ \log P(x_{\epsilon}|y)$, for a small $\epsilon$ such that $\sigma_{\epsilon}\approx 0$. Hence we solve the MAP formulation in Eqs. 14 and 15 in practice. Since, $P(x_{\epsilon}|x_{0})=\mathcal{N}(x_{0},\sigma^{2}_{\epsilon}\mathbb{I})$, for very small values of $\sigma_{\epsilon}$, the distinction between $x_{0}$ and $x_{\epsilon}$ remain insignificant for all practical purposes. The consistency models in  [23] are also learned to map the points on PF ODE trajectory to the corresponding $x_{\epsilon}$ instead of $x_{0}$.
$z^{*}=\arg\max_{z}\ \log P(x_{\epsilon}=f_{\theta}(z,T)|y)$ (14)
$x^{*}_{0}\approx x^{*}_{\epsilon}=f_{\theta}(z^{*},T)$ (15)
Here, we describe in detail the computation of the gradient term $\nabla_{z}\log P(f_{\theta}(z,t)|y)$ in practice. From Eqs. 16 and 17, this requires the estimation of the gradient of log-likelihood i.e. $\nabla_{x_{\epsilon}}\log P(y|x_{\epsilon})$ and the gradient of log-prior i.e. $\nabla_{x_{\epsilon}}\log P(x_{\epsilon})$. The terms $P(y|x_{\epsilon})$ and $P(x_{\epsilon})$ are also referred to as the likelihood and the prior respectively.
$\nabla_{z}\log P(f_{\theta}(z,t)|y)={\left(\frac{\partial f_{\theta}(z,t)}{\partial z}\right)}^{\intercal}\nabla_{x_{\epsilon}}\log P(x_{\epsilon}|y)\bigg{|}_{x_{\epsilon}=f_{\theta}(z,t)}$
$\nabla_{x_{\epsilon}}\log P(x_{\epsilon}|y)\bigg{|}_{x_{\epsilon}=f_{\theta}(z,t)}=\biggl{\{}\nabla_{x_{\epsilon}}\log P(y|x_{\epsilon})+\nabla_{x_{\epsilon}}\log P(x_{\epsilon})\biggr{\}}\bigg{|}_{x_{\epsilon}=f_{\theta}(z,t)}$
Computing the gradient of log-likelihood
$P(y|x_{0})=\mathcal{N}(\mathcal{A}(x_{0}),\sigma^{2}_{y}\mathbb{I})$, and $P(x_{\epsilon}|x_{0})=\mathcal{N}(x_{0},\sigma^{2}_{\epsilon}\mathbb{I})$, given by Eq. 1 and the diffusion perturbation kernel respectively. Since $\sigma_{\epsilon}\approx 0$, $P(x_{0})\approx P(x_{\epsilon})$ and $P(x_{0}|x_{\epsilon})\approx P(x_{\epsilon}|x_{0})$. We can approximate $P(x_{0}|x_{\epsilon})=\mathcal{N}(x_{\epsilon},\sigma^{2}_{\epsilon}\mathbb{I})$, and for a linear forward operator i.e. $\mathcal{A}=H\in\mathbb{R}^{m\times n}$, we can derive using Eq. 7, $P(y|x_{\epsilon})=\mathcal{N}(Hx_{\epsilon},\sigma^{2}_{y}\mathbb{I}+\sigma^{2}_{\epsilon}HH^{\intercal})$. For non-linear $\mathcal{A}$, similar approximations can be made by linearizing it around $x_{\epsilon}$. Given the tractable form of the likelihood term above, the gradient of the log-likelihood is apparent.
Computing the gradient of log-prior
The gradient of the log prior i.e. $\nabla_{x_{\epsilon}}\log P(x_{\epsilon})$ is essentially the score function evaluated at $x_{\epsilon}$. Given a score function $S_{\theta}(x,t)$, learned by the unconditional diffusion model, $\nabla_{x_{\epsilon}}\log P(x_{\epsilon})=S_{\theta}(x_{\epsilon},\epsilon)$. Learning the score function is equivalent to learning a denoiser, and vice-versa [11]. Hereon, we denote the unconditional diffusion model as learning the denoiser $D_{\theta}(x,t)$, from which the score function can be computed as $\nabla_{x_{t}}\log P(x_{t})$ $=\frac{D_{\theta}(x_{t},t)-x_{t}}{\sigma^{2}_{t}}$.
4 Image restoration with MAP-GA
Several image restoration tasks such as inpainting, deblurring, and super-resolution can be modeled as linear inverse problems. For example, in image inpainting [31, 15], given a masked image (and the corresponding binary mask), the goal is to recover (reconstruct) the missing pixels of the masked image. Let $x\in\mathbb{R}^{n}$ denote the original image, $y\in\mathbb{R}^{m}$ denote a masked image with only visible pixels ($m\leq n$) and $H\in\mathbb{R}^{m\times n}$ denotes a corresponding linear forward operator for a given mask. The inpainting problem is characterized by $y=Hx+\eta$, where $\eta\sim\mathcal{N}(0,\sigma^{2}_{y}\mathbb{I})$. Note that for inpainting, a given mask defines $H$ with a defined structure, where the rows of $H$ are one-hot and are orthogonal i.e. $HH^{\intercal}=\mathbb{I}_{m\times m}$. Other image restoration problems such as deblurring and super-resolution can be modeled accordingly with their corresponding forward operators.
We expand Algorithm 1 for image restoration and include all the specific details in Algorithm 2. The core term in the algorithm that needs to be evaluated is $\nabla_{z}\log P(f_{\theta}(z,t)|y)$, which involves the estimation of gradients of the log-likelihood and the log-prior terms (Eqs. 16 and 17). In the algorithm, we make a choice (indicated by the $use\_prior$ keyword) of retaining or dropping the gradient of the log-prior. Dropping the prior term implies a choice of uniform prior, and the algorithm now optimizes for the maximum likelihood estimate instead of the MAP estimate, by considering any sample (consistent with our measurement) on the data manifold to be equally good. This is also the setting considered in the concurrent works  [4, 29].
Note that the algorithm makes use of both the consistency model ($C_{\theta}$) and the denoiser ($D_{\theta}$), which makes our method more demanding compared to other methods that only use the denoiser. We make an argument as follows. We use the pre-trained denoiser and the consistency model from  [11] and  [23] respectively, with noise schedule $\sigma(t)=t$ and $t\in[\epsilon,T]$. Consider the corresponding PF ODE (from Eq. 4) for the above setting, as follows.
$dx=-\frac{D_{\theta}(x,t)-x}{t}dt$
This PF ODE determines the trajectory and solving the trajectory origin $x_{0}$ involves solving the ODE above. Note that this $x_{0}$ is what the consistency model $C_{\theta}$ is trained to predict. As a rough approximation, solving the ODE with a backward Euler discretization step from $t$ to 00, which essentially assumes the trajectory curves are linear, i.e. the Jacobian $\frac{dx}{dt}$ is constant for the interval $[0,t]$, and this gives $C_{\theta}(x_{t},t)=x_{0}\approx D_{\theta}(x_{t},t)$, with the approximation getting more accurate as the trajectory curve gets more linear. In an empirical setting,  [11] also observes trajectories become more linear when $\sigma(t)\to 0$. While more analysis on this is still due, it motivates us to look at the denoiser as a proxy of the consistency model, which gradually becomes more and more accurate as $\sigma(t)\to 0$. Hence, we also consider settings that replace the consistency model with the denoiser in our algorithm.
input :  time schedule: $\tau=[\ \tau_{n},\tau_{n-1},..\tau_{1},\tau_{0}\ ]$,
noise schedule: $\sigma(.)$, denoiser: $D_{\theta},$
consistency model: $C_{\theta}$,
measurement: $y$, learning rate: $\lambda,$
num gradient ascent iter: num_iter,
boolean: use_prior (default True),
forward operator matrix: H,
measurement noise: $\sigma_{y}$
(Note: $\tau_{0}=\epsilon$, $\tau_{n}=T$),
$z\sim\mathcal{N}(0,\sigma^{2}_{\tau_{n}}\mathbb{I})$
[…]
       $t=\tau_{i}$
[…]
             $\hat{x}_{\epsilon}=C_{\theta}(z,t)$
             $grad_{likelihood}=\frac{H^{\intercal}(\frac{\sigma^{2}_{y}}{\sigma^{2}_{\epsilon}}\mathbb{I}+HH^{\intercal})^{-1}(y-H\hat{x}_{\epsilon})}{\sigma^{2}_{\epsilon}}$
             if use_prior then
                  $grad_{prior}=\frac{D_{\theta}(\hat{x}_{\epsilon},\epsilon)-\hat{x}_{\epsilon}}{\sigma^{2}_{\epsilon}}$
             end if
            else
                  $grad_{prior}=0$
             end if
            $grad_{posterior}=grad_{likelihood}+grad_{prior}$
             $grad=\left(\frac{\partial C_{\theta}(z,t)}{\partial z}\right)^{\intercal}grad_{posterior}$
             $z=z+\lambda*grad$
       end for
      $\hat{x}_{\epsilon}=C_{\theta}(z,t)$
       $z=\mathcal{N}(\hat{x}_{\epsilon},\sigma^{2}_{\tau_{i-1}}-\sigma^{2}_{\tau_{0}}\mathbb{I})$
end for
output : z
Algorithm 2 MAP-GA for Image restoration
When using Algorithm 2 for noisy image restoration ($\sigma_{y}>0$) in practice, we observe that it requires careful tuning of the learning rate and other hyperparameters. To avoid such additional and sensitive hyperparameters, we present Algorithm 3 for noisy image restoration, based on our empirical observations in Tab. 3. The motivation behind Algorithm 3, is to find an approximate solution using MAP-GA at diffusion time $t=\tau$ where $\sigma_{\tau}=\sigma_{y}$, and then use it as an initialization for PGDM [21] for $\sigma_{t}<\sigma_{y}$. More specifically, we use MAP-GA until $\sigma_{t}=\sigma_{y}$, to estimate $x_{\sigma_{y}}$ and later use this as an initialization to PGDM for $\sigma_{t}<\sigma_{y}$. We do not use the prior term for the MAP-GA part in Algorithm 3, as we find it more effective.
input :  noise schedule: $\sigma(.)$, denoiser: $D_{\theta},$
consistency model: $C_{\theta}$,
measurement: $y$, learning rate: $\lambda,$
num gradient ascent iter: num_iter,
forward operator matrix: H,
measurement noise: $\sigma_{y}$
$\tau^{map-ga}=[\tau^{map}_{n},\tau^{map}_{n-1},..\tau^{map}_{1},\tau^{map}_{0}]$,
$\tau^{pgdm}=[\tau^{pgdm}_{m},\tau^{pgdm}_{m-1},..\tau^{pgdm}_{1},\tau^{pgdm}_{0}]$,
(Note: $\sigma_{\tau^{map}_{0}}=\sigma_{y}$ and $\tau^{map}_{n}=T$)
(Note: $\sigma_{\tau^{pgdm}_{m}}=\sigma_{y}$ and $\tau^{pgdm}_{0}=\epsilon$),
=======MAP-GA=======
$z\sim\mathcal{N}(0,\sigma^{2}_{\tau^{map}_{n}}\mathbb{I})$
[…]
       $t=\tau^{map}_{i}$
[…]
             $\hat{x}_{\epsilon}=C_{\theta}(z,t)$
             $grad_{likelihood}=\frac{H^{\intercal}(\frac{\sigma^{2}_{y}}{\sigma^{2}_{\epsilon}}\mathbb{I}+HH^{\intercal})^{-1}(y-H\hat{x}_{\epsilon})}{\sigma^{2}_{\epsilon}}$
             $grad_{posterior}=grad_{likelihood}$
             $grad=\left(\frac{\partial C_{\theta}(z,t)}{\partial z}\right)^{\intercal}grad_{posterior}$
             $z=z+\lambda*grad$
       end for
      $\hat{x}_{\epsilon}=C_{\theta}(z,t)$
       $z=\mathcal{N}(\hat{x}_{\epsilon},\sigma^{2}_{\tau^{map}_{i-1}}-\sigma^{2}_{\tau^{map}_{0}}\mathbb{I})$
end for
========PGDM========
$x_{\tau^{pgdm}_{m}}=z$
[…]
       $t=\tau^{pgdm}_{i}$
       $\hat{x}_{t}=D_{\theta}(z,t)$
       $\mu_{t}=\hat{x}_{t}+\sigma^{2}_{t}*\nabla_{x_{t}}\log P(y|x_{t})$pgdm update*
       $x_{\tau^{pgdm}_{i-1}}=\mathcal{N}(\mu_{t},\sigma^{2}_{\tau^{pgdm}_{i-1}}-\sigma^{2}_{\tau^{pgdm}_{0}}\mathbb{I})$
end for
output : $x_{\tau^{pgdm}_{0}}$
Algorithm 3 MAP-GA-PGDM Image restoration
box50 half expand box25 sr2x altlines deblur supres4x
[…]
MAP-GA(D,NP) 32.312 0.096 37.986 0.157 $93.068$ $0.419$ 11.019 0.019 30.072 0.034 18.282 0.015 19.704 0.007 $66.968$ $0.148$
MAP-GA(D) $35.761$ $0.100$ $43.315$ $0.164$ $106.32$ $0.430$ $11.169$ $0.019$ $33.188$ $0.036$ $19.103$ $0.016$ $19.779$ $0.008$ $78.791$ $0.179$
MAP-GA(NP) $34.944$ $0.113$ $39.243$ $0.162$ 69.004 0.388 $12.818$ $0.027$ $30.303$ $0.035$ $20.321$ $0.018$ $22.090$ $0.008$ 46.349 0.112
MAP-GA $36.733$ $0.113$ $41.151$ $0.164$ $87.952$ $0.400$ $12.836$ $0.027$ $33.752$ $0.036$ $20.895$ $0.018$ $21.314$ $0.010$ $59.624$ $0.120$
PGDM [21] $49.370$ $0.151$ $54.261$ $0.245$ $127.95$ $0.479$ $14.255$ $0.021$ $38.433$ $0.046$ $20.446$ $0.019$ $19.857$ $0.007$ $89.614$ $0.238$
DDRM [12] $51.477$ $0.165$ $56.643$ $0.264$ $136.06$ $0.492$ $15.000$ $0.023$ $35.033$ $0.041$ $19.331$ $0.017$ $23.195$ $0.009$ $78.712$ $0.235$
CT-ZSIE [23] $38.017$ $0.129$ $44.152$ $0.191$ $70.634$ $0.424$ $13.040$ $0.025$ $42.500$ $0.060$ $26.737$ $0.029$ $29.223$ $0.013$ $56.698$ $0.116$
Table 1: Noiseless image restoration on ImageNet64 1K validation set using MAP-GA and variants. The setting MAP-GA denotes the default Algorithm 2, MAP-GA(NP) denotes MAP-GA with no prior, MAP-GA(D) denotes MAP-GA with denoiser replacing the consistency model, and MAP-GA(D,NP) denotes MAP-GA with no prior and the denoiser replacing the consistency model.
box50 half expand box25 sr2x altlines deblur supres4x
[…]
MAP-GA(D,NP) 70.880 0.128 68.022 0.269 $174.96$ $0.695$ $18.221$ $0.028$ 22.819 $0.059$ $13.632$ $0.036$ $81.461$ $0.214$ $59.895$ $0.205$
MAP-GA(D) $74.995$ $0.137$ $75.748$ $0.280$ $197.91$ $0.696$ 16.795 0.025 $24.420$ $0.063$ $13.669$ $0.033$ $86.585$ $0.220$ $63.308$ $0.214$
MAP-GA(NP) $98.749$ $0.165$ $86.975$ $0.297$ 155.55 0.645 $30.788$ $0.046$ $24.434$ 0.054 $16.126$ $0.034$ 80.337 0.211 49.622 0.146
MAP-GA $108.87$ $0.171$ $85.936$ $0.291$ $175.27$ $0.652$ $34.711$ $0.053$ $25.666$ $0.056$ $15.139$ $0.036$ $84.971$ $0.214$ $68.156$ $0.169$
PGDM [21] $120.82$ $0.194$ $94.920$ $0.360$ $227.97$ $0.765$ $28.357$ $0.041$ $27.927$ $0.070$ $14.211$ $0.037$ $94.629$ $0.227$ $77.533$ $0.248$
DDRM [12] $131.86$ $0.198$ $101.86$ $0.379$ $224.93$ $0.778$ $29.571$ $0.041$ $23.686$ $0.065$ 12.040 0.026 $105.28$ $0.266$ $75.923$ $0.251$
CT-ZSIE [23] $118.27$ $0.209$ $121.11$ $0.375$ $200.03$ $0.704$ $34.246$ $0.046$ $47.353$ $0.145$ $24.663$ $0.070$ $97.343$ $0.264$ $50.608$ $0.157$
Table 2: Noiseless image restoration on $100$ LSUNCat256 images using MAP-GA and variants. The setting MAP-GA denotes the default Algorithm 2, MAP-GA(NP) denotes MAP-GA with no prior, MAP-GA(D) denotes MAP-GA with denoiser replacing the consistency model, and MAP-GA(D,NP) denotes MAP-GA with no prior and the denoiser replacing the consistency model.
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Figure 1: Left to right: original image, and mask settings: box50, half, expand, box25, sr2x, altlines
5 Experiments
We consider the tasks of image inpainting, deblurring, and 4×4$\times$4 × super-resolution. Inspired by  [13], for image inpainting, we evaluate the performance across six different mask settings (Fig. 1) denoting varying levels of degradation. The mask settings box50 and box25 indicate a square crop at the center of the image, with the crop width equal to 50% and 25% of the image width respectively. In half, we mask out the right half of the image, expand is the complement of box25. sr2x denotes a 2×2$\times$2 × super-resolution mask, and altlines masks out alternate rows of pixels. We evaluate the restored images using FID and LPIPS metrics.
We evaluate the performance of MAP-GA (Algorithm 2) on ImageNet [18] 1K validation set with $64\times 64$ resolution (in Tab. 1) and on $100$ random images of LSUNCat [32] with $256\times 256$ resolution (in Tab. 2). We use the pre-trained denoisers and the consistency models from  [11, 23] with their default settings. In the experiments, the setting MAP-GA denotes the default Algorithm 2, MAP-GA(NP) denotes MAP-GA with no prior, MAP-GA(D) denotes MAP-GA with denoiser replacing the consistency model, and MAP-GA(D,NP) denotes MAP-GA with no prior and the denoiser replacing the consistency model. We compare against DDRM [12] and PGDM [21] (both only use the denoiser) and against the zero-shot image editing (CT-ZSIE) algorithm proposed in  [23] which only use the consistency model. The results from Tabs. 1 and 2 show MAP-GA and variants outperform DDRM, PGDM, and CT-ZSIE with a significant margin on several tasks.
Note that MAP-GA uses gradient ascent (first-order gradient-based optimization algorithm) to optimize the underlying MAP objective which is typically highly non-convex and is not guaranteed to find the global optima. However, it could converge with an initialization closer to the global optima. To corroborate this, we design the following toy experiment. We consider the noiseless inpainting task from earlier, but, in Algorithm 2, we set $\tau_{n}=\hat{\tau}\ll T$, instead of the default setting $\tau_{n}=T$ and we also initialize $z\sim\mathcal{N}(x_{0},\sigma^{2}_{\hat{\tau}}\mathbb{I})$, where $x_{0}$ is the corresponding ground truth image for the measurement $y$. From Tab. 3, we observe significant improvements in the performance. (We use $\hat{\tau}=0.5$, note that $\sigma_{0.5}=0.5$, as we use the noise schedule $\sigma_{t}=t$). This shows that MAP-GA is only limited by the choice of optimizer and reinforces the need for better optimization algorithms. While it is important to consider better design choices (for the schedules and hyperparameters), adaptive-gradient-based optimizers (such as momentum, Adam), or higher-order methods, we leave this for future work as it requires a thorough analysis.
In Sec. 5, we report the results on noisy inpainting using Algorithm 3 on the ImageNet64 1K validation set. In Algorithm 3, and all its variants, we fix $m=20$ time steps for PGDM. Even with high levels of measurement noise, MAP-GA-PGDM shows promising improvements over PGDM. In all our experiments, for MAP-GA and variants, we fix a budget of 1000 steps and run ablations for $(num\_steps,num\_iter)$ from the set {(20,50),(50,20),(100,10),(200,5),(250,4),(500,2),(1000,1)}. For DDRM, PGDM, and CT-ZSIE, we run ablations for $num\_steps$ from the set {20,50,100,200,250,500,1000}. The learning rate in all our experiments was set to $\sigma^{2}_{y}+\sigma^{2}_{\epsilon}$.
box50 half expand box25 sr2x altlines
[…]
MAP-GA(D,NP) $26.661$ $0.043$ $30.621$ $0.064$ $73.510$ $0.155$ 9.165 $0.010$ $29.878$ $0.027$ 16.971 0.012
MAP-GA(D) $34.689$ $0.052$ $41.098$ $0.083$ $90.768$ $0.175$ $10.891$ $0.011$ $34.334$ $0.031$ $18.718$ $0.013$
MAP-GA(NP) 25.205 0.043 27.208 0.047 42.671 0.075 $10.936$ $0.017$ 27.496 0.023 $18.732$ $0.013$
MAP-GA $29.508$ $0.045$ $33.653$ $0.050$ $73.318$ $0.106$ $11.257$ $0.016$ $34.364$ $0.026$ $20.041$ $0.013$
PGDM [21] $31.952$ $0.056$ $36.639$ $0.082$ $71.352$ $0.138$ $10.886$ 0.009 $33.199$ $0.041$ $19.551$ $0.018$
Table 3: Noiseless inpainting on ImageNet64 1K validation set. Using the ground truth image ($x_{0}$) for the measurement $y$, we create a sample at $t=0.5$ via ($x_{0.5}=x_{0}+0.5*\eta$, where, $\eta\sim\mathcal{N}(0,\mathbb{I}))$ and initialize Algorithm 2 with $z=x_{0.5}$, and $\tau_{n}=0.5$
box50 half expand box25 sr2x altlines
[…]
MAP-GA-PGDM(D) $0.05$ 56.173 0.125 $62.026$ $0.193$ $108.16$ $0.438$ 38.725 0.039 $57.41$ $0.080$ $46.046$ $0.046$
MAP-GA-PGDM $0.05$ $58.283$ $0.147$ 61.588 0.184 91.508 0.406 $44.667$ $0.085$ 57.308 0.073 45.265 0.042
PGDM [21] $0.05$ $77.824$ $0.175$ $80.248$ $0.257$ $135.99$ $0.495$ $53.136$ $0.049$ $86.289$ $0.126$ $66.203$ $0.066$
[…]
MAP-GA-PGDM(D) $0.1$ 72.543 0.166 $79.535$ $0.230$ $114.51$ $0.464$ 57.556 0.076 $76.733$ $0.145$ $65.154$ $0.096$
MAP-GA-PGDM $0.1$ $74.130$ $0.191$ 78.322 0.225 103.85 0.440 $65.720$ $0.161$ 76.248 0.134 63.134 0.089
PGDM [21] $0.1$ $96.485$ $0.216$ $99.170$ $0.286$ $145.40$ $0.519$ $78.620$ $0.100$ $109.47$ $0.231$ $90.925$ $0.138$
[…]
Table 4: Noisy inpainting on ImageNet64 1K validation set. $\sigma_{y}$ denotes the measurement noise. The setting MAP-GA-PGDM denotes the default Algorithm 3, MAP-GA-PGDM(D) denote MAP-GA-PGDM with denoiser replacing the consistency model.
6 Discussion
6.1 Runtime comparison
Tab. 5 compares the wall-clock time of MAP-GA and variants against DDRM, PGDM, and CT-ZSIE for image inpainting on ImageNet64. To ensure a fair comparison, we keep the batch size fixed at 50 for all the methods and compare their runtime per iteration. For DDRM, CT-ZSIE, and PGDM, it is the total runtime divided by $num\_steps$ (i.e. reverse diffusion time steps), while for MAP-GA and variants, it is the effective runtime per $num\_steps$ per $num\_iter$ (i.e. it is the runtime when $num\_steps=num\_iter=1$). MAP-GA and variants are 1.5×1.5$\times$1.5 × to 2×2$\times$2 × slower per iteration than PGDM and 3×3$\times$3 × to 4×4$\times$4 × slower than DDRM and CT-ZSIE.
Method Runtime per iteration
DDRM [12] 150 ms
CT-ZSIE [23] 150 ms
PGDM [21] 304 ms
MAP-GA(D,NP) 456 ms
MAP-GA(D) 602 ms
MAP-GA(NP) 455 ms
MAP-GA 603 ms
Table 5: Runtime comparison on NVIDIA A100 40GB GPU
6.2 Concurrent works
The algorithms presented in this paper are similar to ZSIR [4], and DMPlug [29] from a practical perspective. However, our MAP formulation is novel, has a strong theoretical motivation, and connects the PF ODE and the consistency model with the MAP optimization for solving inverse problems. Unlike ZSIR and DMPlug, we consider the prior term and show that the gradient of the log-prior is tractable, making the gradient of the log-posterior tractable. We show that MAP-GA is only limited by the optimizer’s choice in practice. ZSIR and DMPlug approximate the PF ODE trajectory origin $x_{0}$ with a multi-step approximation using the denoiser and hence require backpropagation through the chain of cascaded functions to optimize for the loss objective. In our case, all MAP-GA variants require a single vector-Jacobian product per iteration.
Figure 2: Noiseless inpainting task. Left to right: original image, masked image, MAP-GA(Algorithm 2), PGDM($\sigma_{y}=0$). Top to bottom: half mask, expand mask, box50 mask
Figure 3: Noisy inpainting task. Left to right: original image, masked image, MAP-GA-PGDM(Algorithm 3, $\sigma_{y}=0.1$), PGDM($\sigma_{y}=0.1$). Top to bottom: sr2x mask, box25 mask
7 Conclusion
In this paper, we proposed a novel MAP formulation for solving inverse problems using pre-trained unconditional diffusion models. Note that conditional generation is a core requirement in solving inverse problems. We connect the Probability Flow ODE and the consistency model with the optimization process for the MAP objective in tasks that involve conditional generation. We show that the gradient of the MAP objective is tractable, allowing the use of gradient-based optimization methods. To use our framework in practice, we propose an algorithm with a multi-step gradient ascent strategy for MAP optimization. We propose practically effective algorithms for image restoration using our framework (extendable to other general inverse problems). We validate our algorithms with extensive experiments on image deblurring, super-resolution, and inpainting.
Acknowledgements
This work is part of the Marie Skłodowska-Curie Actions project MODELAIR, funded by the European Commission under the Horizon Europe program through grant agreement no. 101072559. The computations and the data handling were enabled by resources provided by the National Academic Infrastructure for Supercomputing in Sweden (NAISS), partially funded by the Swedish Research Council through grant agreement no. 2022-06725. Bharath thanks Sebastian Gerard and Heng Fang for their feedback on improving the presentation of this paper.